# Eccentric and cool: a high-spectral resolution view of

### v1

@@
Eccentric and cool: a high-spectral resolution view of
@@
(Received 05 January 2026; accepted 31 January 2026)
Discovered almost 10 years ago, the giant planet 51 Eridani b is one of the least separated ($\approx$0.2′′) and faintest (J $\approx$ 19.74 mag) directly imaged exoplanets known to date. Its atmospheric properties have been thoroughly investigated through low- and medium-resolution spectroscopic observations, enabling robust characterization of the planet’s bulk parameters. However, the planet’s intrinsically high contrast renders high-resolution spectroscopic observations difficult, despite their potential to yield key measurements essential for a more comprehensive characterization. This study seeks to constrain the planet’s radial velocity, enabling a full 3D orbital solution when integrated with previous measurements. We have obtained 4 high-contrast high-resolution (R $\approx$ 140 000) spectroscopic datasets of the planet, collected over a two-year interval with the HiRISE visitor instrument at the VLT to derive the planet’s radial velocity. Using self-consistent models of atmosphere, we were able to derive the radial velocity of the planet at each of the 4 epochs. These radial velocity measurements were then used in combination with all existing relative astrometry in order to constrain the orbit of the planet.Our radial velocity measurements allow us to break the degeneracy along the line of sight, making it now possible the unambiguous interpretation of the phase curve of the companion. We further constrain the orbital parameters, particularly the eccentricity, for which we derive e = $0.55_{-0.07}^{+0.03}$. The relatively high eccentricity indicates that the system has experienced dynamical interactions induced by an external perturber. We place constraints on the mass and semi-major axis of a hypothetical, unseen outer planet capable of producing the observed high eccentricities.
@@
Instrumentation: high angular resolution – Instrumentation: adaptive optics – Instrumentation: spectrographs – Techniques: high-angular resolution – Techniques: spectroscopy – Infrared: planetary systems
@@ 1 Introduction
Eccentricity, for instance, is an efficient indicator of planetary formation and dynamical evolution. While planets forming in protoplanetary disks are expected to form on circular orbits, they can develop non-zero eccentricities through dynamical processes such as planet–planet scattering (Ford and Rasio 2008), Kozai–Lidov oscillations (Kozai 1962; Lidov 1962), planet–disk interactions (Li and Lai 2023), or stellar flybys (Pfalzner et al. 2024). A population-level dichotomy exists between brown dwarf and giant planet companions, with brown dwarfs peaking at high eccentricities ($e\approx 0.6$$0.9$) and giant planets typically exhibiting low eccentricities (Bowler et al. 2020). The giant planet 51 Eridani b (hereafter 51 Eri b) is an exception, as recent studies suggest a potentially high eccentricity (Rosa et al. 2020; Dupuy et al. 2021; Balmer et al. 2025), indicative of past dynamical interactions.
Discovered in 2014 by the Gemini Planet Imager Exoplanet Survey (Macintosh et al. 2015), 51 Eri b is orbiting an F0 star that is a member of the $\beta$ Pictoris moving group ($24\pm 3$ Myr; Bell et al. 2015). Updated stellar parameters yield $R_{\star}=1.45\pm 0.02,R_{\odot}$ and $M_{\star}=1.550\pm 0.006,M_{\odot}$, with an age of $23\pm 1$ Myr (Elliott et al. 2024). The system also hosts the wide, co-moving M-dwarf binary GJ 3305 AB at a projected separation of $\sim$2000,au (Feigelson et al. 2006; Kasper et al. 2007).
@@
In this work, we present nearly two years of high-resolution ($R=140,000$) direct spectroscopic observations of 51 Eri b obtained with the HiRISE instrument (Vigan et al. 2024), combining SPHERE (Beuzit et al. 2019) and CRIRES (Dorn et al. 2023). We describe the observations and data reduction in Sec. 2, present the radial velocity monitoring in Sec. 3, update the orbital solution by combining RVs with literature astrometry in Sec. 4, and summarize our conclusions in Sec. 5.
@@ 2 Observations and data reduction
We observed the 51 Eri system at four epochs with VLT/HiRISE (see Table 3), following the same observing strategy as for AF Lep b described in Denis et al. (2025). For the first three epochs, sky backgrounds were acquired after the science exposures to subtract the leakage associated with the HiRISE MACAO guide fiber (Vigan et al. 2024). After a technical intervention in January 2025, the leakage issue was fully resolved, and subsequent observations relied on internal daytime backgrounds.
@@
Standard CRIRES calibrations were acquired automatically the next morning based on the science observations of each night. This includes dark, flat fields and wavelength calibration files, as detailed in the CRIRES calibration plan. HiRISE observations do not require any specific internal calibrations in addition to standard daily calibrations of the spectrograph.
Data reduction was performed using the public HiRISE pipeline (Costes et al. 2024), which applies background and flat-field corrections, combines exposures, extracts the spectra, recalibrates the wavelength, and corrects for barycentric motion, while providing calibrated stellar and companion spectra22footnotemark: 2, sky and instrumental response, and associated uncertainties. The pipeline is described in detail in Denis et al. (2025) and now includes an optimal extraction scheme (Horne 1986), representing a significant improvement over earlier reductions.
@@ 3.1 Forward modeling analysis
To measure the radial velocity (RV) of the companion, we compared the data to self-consistent atmospheric models using the forward-modeling code ForMoSA (Petrus et al. 2021). ForMoSA operates within a Bayesian framework and allows the joint analysis of spectroscopic and photometric data over a wide range of wavelengths and resolutions. Atmospheric model grids spanning effective temperature, gravity, metallicity, and related parameters are interpolated at each iteration, with optional modifications such as Doppler shifts and rotational broadening applied before computing the likelihood. The output consists of the best-fitting atmospheric parameters and their uncertainties.
We adopted the radiative–convective equilibrium model Exo-REM (Charnay et al. 2018) recomputed at a spectral resolution of 200 000, using low-resolution Exo-REM volume mixing ratio profiles in tandem with Exo_k. This model is referred to as Exo-REM/Exo_k (Radcliffe et al., in prep). A cloud-free model was used, with priors informed by Balmer et al. (2025). Given the high spectral resolution and narrow wavelength coverage, the data are expected to probe the upper atmospheric layers above the cloud base, as discussed in Xuan et al. (2022) and Denis et al. (2025). To compare the model to the data, we followed the framework described in Denis et al. (2025).
The results are summarized in Table 1, which reports the planetary RV corrected for barycentric motion and the stellar RV ($12.6\pm 0.3$$\mathrm{km}\,\mathrm{s}^{-1}$; Gontcharov 2006), as well as the planetary $v\sin i$ . The $v\sin i$ measured at the fourth epoch is inconsistent with the other epochs, likely due to instrumental fringing (see Appendix B). We therefore rely on the first three epochs to estimate the rotational broadening and derive $\hbox{$v\sin i$}=9.9\pm 2.3$$\mathrm{km}\,\mathrm{s}^{-1}$, keeping in mind that the fringing effect still affect these data. Assuming $\sin i=1$ and adopting the planetary radius $R_{p}=1.3\pm 0.03,R_{\mathrm{Jup}}$ (Balmer et al. 2025), we can derive an upper limit on the rotation period of the planet ($\sim 16.4\pm 3.8$ h). We thus estimate $P_{\mathrm{rot}}\lesssim 24$ h at 2 $\sigma$.
Table 1: Results on 51 Eri b for the 4 nights of observation333
@@ 3.2 Cross-correlation function
To assess the presence of individual molecules in the atmosphere, it is convenient to use cross-correlation function (CCF) analysis, where we correlate molecular templates with the data. In the H band, at the temperature of 51 Eri b, the dominant species are H2O, CH4 and CO. To generate the templates for these molecules, we used the pressure-temperature profiles and the volume mixing ratio profiles of Exo-REM in tandem with Exo_K. For the derivation of the CCF, we refer to Appendix C.
There appears to be no correlation between the CCF and the derived RV uncertainty. For the first and third epochs, where detection confidence is highest, the derived RV uncertainties are more than twice those of the final epoch. Although it is natural to expect RV uncertainty to correlate with CCF S/N, we are in a regime where resolution is also an important parameter. The lower vsini of the last epoch compared to the other epochs keeps the effective resolution of the template higher (see Table 1). This results in sharper variations of the log likelihood near the RV of the planet and therefore less uncertainty in its determination.
Figure 1 shows the combined CCF from the 4 epochs. Before combining the 4 CCFs, we shifted each individual CCF to the restframe of the planet to account for the variations in RV across the different epochs. In addition to the individual molecular templates, we also considered a full template including H2O + CH4 + CO. In general, the molecular detection at individual epochs are marginal, but the combined CCF exhibits a reliable detection of H2O and CH4, and a tentative detection of CO, which confirms the recent findings of Madurowicz et al. (2025).
Figure 1: Combined CCF from the 4 nights. The combined CCF has been computed by shifting the each individual CCF to a RV of 0 $\mathrm{km}\,\mathrm{s}^{-1}$ in order to account for the variations of the RV of the companion across the different epochs.
@@ 4 Updated orbital solution
Table 2: Orbital parameters inferred in this work.444\$a\$\$a\$footnotetext: The periastron passage is computed as $\tau=\frac{t_{\mathrm{p}}-t_{\mathrm{ref}}}{P}$ where tref is the reference time at epoch J2020.0 ($t_{\mathrm{ref}}$ = 58849 MJD) and tp is the time of periastron.
@@
To update the orbit of 51 Eri b, we compiled relative astrometric data from Gemini/GPI and Keck/NIRC2 (Macintosh et al. 2015; Rosa et al. 2020), VLT/SPHERE (Maire et al. 2019) and JWST/NIRCam (Balmer et al. 2025). We used the Markov Chain Monte Carlo (MCMC) framework designed by Beust et al. (2016) that makes use of universal Keplerian variables with Stumpff functions (Danby 1987). This formulation achieves more efficient convergence in the case of astrometric orbits sampled over a small orbital arc, as in the present case. The MCMC was run using 8 walkers. The burn-in and convergence was assessed using Gelman-Rubin statistics (Gelman et al. 2003; Ford 2006). Before beginning the inference of the chains, we require that $\hat{R}(z)<1.1$ and $\hat{T}(z)>100$ (see Sects. 3.3 and 3.4 of Ford 2006). Then, we let the chains be filled until $\hat{R}(z)<1.01$ and $\hat{T}(z)>1000$. The filled chains constitutes the output of the orbital fit.
@@
We do not directly fit for classical orbital elements: the semi-major axis $a$, the eccentricity $e$, the inclination $i$, the argument of periastron $\omega$, the longitude of ascending node $\Omega$, and the time of periastron passage $t_{p}$. Ford (2006) suggests that working with combinations of these elements allows an optimization of the Markov Chains. We specifically fit for $\ln q$, $e\cos(\Omega+\omega)$, $e\sin(\Omega+\omega)$, $sin(i/2)\cos\Omega$, $\sin(i/2)\sin\Omega$, $\ln P$, and $s$. Here $q$ is the periastron, $P_{q}$ is the orbital period of a circular orbit with radius $q$, and $s$ is the universal variable (a reformulation of the eccentric anomaly, Danby 1987) at a reference epoch. Note that $P$ can also be defined as $P_{q}=2\pi qv_{q}$, where $v_{q}$ is the velocity at periastron. From any solution with those fitted parameters, the classical elements can be derived unambiguously.
@@
The updated orbit and associated parameters are presented in Fig. 2 and Table 2. The full posterior distributions of orbital parameters are presented in Fig. 6. The left panel of Fig. 2 shows the predicted RV differences between the planet and the star as a function of time. The middle panel is a zoomed-in version centered on the HiRISE measurements and the right panel displays the astrometric orbit of the companion. These plots are generated using 500 orbits randomly drawn from the posterior distributions.
@@
The inclusion of our RV measurements rules out a significant number of orbital solutions, as can be seen on the left panel of Fig. 2. The middle panel shows that a future RV measurement of the planet at the end of 2026 will potentially completely reject the last population of low-eccentricity orbits and place even tighter constraints on the eccentricity. As illustrated in Fig. 6, adding the HiRISE RV measurements significantly improves the orbit. We now estimate the eccentricity to lie within the range 0.2–0.7 at $5\sigma$, with a median at $e=0.55$, completely ruling out very eccentric orbits, above 0.7.
@@
The HiRISE measurements also allow to break the typical degeneracy arising with the argument of periastron, $\omega$, and the longitude of ascending node, $\Omega$, when using only astrometric data to fit orbits of exoplanets. This is because the computation of the astrometry of the planet from the orbital parameters is invariant to the transformation ($\omega^{\prime}=\omega+\pi$, $\Omega^{\prime}=\Omega+\pi$), typically producing double-peaked structures in the posterior distributions of these parameters. As a consequence, including our RV measurements enable the determination of the direction of the motion of the companion along the line of sight. This allows us to compute a phase curve of the companion (see the right panel of Fig. 7), which is essential for planning follow-up observations using reflected-light imaging or spectroscopy with future instrumentation. This is particularly pertinent in this case, since the reflected-light contrast is predicted to be higher than the thermal-light contrast for this planet, mostly under favourable phase functions (see Appendix. E).

### v2

# Eccentric and cool: A high-spectral-resolution view of

@@
Eccentric and cool: A high-spectral-resolution view of
@@
(Received 05 January 2025; accepted 31 January 2026)
Discovered almost ten years ago, the giant planet 51 Eridani b is one of the least separated ($\approx$0.2′′) and faintest (J $\approx$ 19.74 mag) directly imaged exoplanets known to date. Its atmospheric properties have been thoroughly investigated through low- and medium-resolution spectroscopic observations, enabling the robust characterization of the planet’s bulk parameters. However, the planet’s intrinsically high contrast renders high-resolution spectroscopic observations difficult, despite their potential to yield key measurements essential for a more comprehensive characterization. This study sought to constrain the planet’s radial velocity, enabling a full 3D orbital solution when integrated with previous measurements. We obtained four high-contrast, high-resolution (R $\approx$ 140 000) spectroscopic datasets of the planet, collected over a two-year interval with the HiRISE visitor instrument at the VLT to derive the planet’s radial velocity. Using self-consistent models of atmosphere, we were able to derive the radial velocity of the planet at each of the four epochs. These radial-velocity measurements were then used in combination with all existing relative astrometry in order to constrain the orbit of the planet.Our radial velocity measurements allowed us to break the degeneracy along the line of sight, making the unambiguous interpretation of the phase curve of the companion possible. We further constrained the orbital parameters, and particularly the eccentricity, for which we derive e = $0.55_{-0.07}^{+0.03}$. The relatively high eccentricity indicates that the system has experienced dynamical interactions induced by an external perturber. We place constraints on the mass and semimajor axis of a hypothetical, unseen outer planet capable of producing the observed high eccentricities.
@@
Instrumentation: high angular resolution – Instrumentation: adaptive optics – Instrumentation: spectrographs – Techniques: high angular resolution – Techniques: spectroscopy – Infrared: planetary systems
@@ 1 Introduction
Eccentricity, for instance, is an efficient indicator of planetary formation and dynamical evolution. While planets forming in protoplanetary disks are expected to form on circular orbits, they can develop nonzero eccentricities through dynamical processes such as planet–planet scattering (Ford and Rasio 2008), Kozai–Lidov oscillations (Kozai 1962; Lidov 1962), planet–disk interactions (Li and Lai 2023), or stellar flybys (Pfalzner et al. 2024). A population-level dichotomy exists between brown-dwarf and giant-planet companions, with brown dwarfs peaking at high eccentricities ($e\approx 0.6$$0.9$) and giant planets typically exhibiting low eccentricities (Bowler et al. 2020). The giant planet 51 Eridani b (hereafter 51 Eri b) is an exception, as recent studies suggested a potentially high eccentricity (Rosa et al. 2020; Dupuy et al. 2021; Balmer et al. 2025), which is indicative of past dynamical interactions.
Discovered in 2014 by the Gemini Planet Imager Exoplanet Survey (Macintosh et al. 2015), 51 Eri b is orbiting an F0 star that is a member of the $\beta$ Pictoris moving group ($24\pm 3$ Myr; Bell et al. 2015). Updated stellar parameters yield $R_{\star}=1.45\pm 0.02,R_{\odot}$ and $M_{\star}=1.550\pm 0.006,M_{\odot}$, with an age of $23\pm 1$ Myr (Elliott et al. 2024). The system also hosts the wide, co-moving M-dwarf binary GJ 3305 AB at a projected separation of $\sim$2000 au (Feigelson et al. 2006; Kasper et al. 2007).
@@
In this paper, we present nearly two years of high-resolution ($R=140,000$) direct spectroscopic observations of 51 Eri b obtained with the HiRISE instrument (Vigan et al. 2024), which combines SPHERE (Beuzit et al. 2019) and CRIRES (Dorn et al. 2023). We describe the observations and data reduction in Sect. 2 and present the radial-velocity monitoring in Sect. 3. In Sect. 4, we provide an update of the orbital solution realized by combining RVs with literature astrometry. We summarize our conclusions in Sect. 5.
@@ 2 Observations and data reduction
We observed the 51 Eri system at four epochs with VLT/HiRISE (see Table 3) following the same observing strategy for AF Lep b described in Denis et al. (2025). For the first three epochs, sky backgrounds were acquired after the science exposures to subtract the leakage associated with the HiRISE MACAO guide fiber (Vigan et al. 2024). After a technical intervention in January 2025, the leakage issue was fully resolved, and subsequent observations relied on internal daytime backgrounds.
@@
Standard CRIRES calibrations were acquired automatically the next morning based on the science observations of each night. These include dark, flat fields and wavelength-calibration files, as detailed in the CRIRES calibration plan. HiRISE observations do not require any specific internal calibrations in addition to standard daily calibrations of the spectrograph.
Data reduction was performed using the public HiRISE pipeline (Costes et al. 2024), which applies background and flat-field corrections, combines exposures, extracts the spectra, recalibrates the wavelength, and corrects for barycentric motion, while providing calibrated stellar and companion spectra,22footnotemark: 2sky and instrumental response, and associated uncertainties. The pipeline is described in detail in Denis et al. (2025) and now includes an optimal extraction scheme (Horne 1986), representing a significant improvement over earlier reductions.
@@ 3.1 Forward modeling analysis
To measure the radial velocity (RV) of the companion, we compared the data to self-consistent atmospheric models using the forward-modeling code ForMoSA (Petrus et al. 2021). ForMoSA operates within a Bayesian framework and allows the joint analysis of spectroscopic and photometric data over a wide range of wavelengths and resolutions. Atmospheric model grids spanning effective temperature, gravity, metallicity, and related parameters are interpolated at each iteration, with optional modifications such as Doppler shifts and rotational broadening to be applied before computing the likelihood. The output consists of the best-fitting atmospheric parameters and their uncertainties.
We adopted the radiative–convective equilibrium model Exo-REM (Charnay et al. 2018) recomputed at a spectral resolution of 200 000, using low-resolution Exo-REM volume mixing ratio profiles in tandem with Exo_k. This model is referred to as Exo-REM k26 (Radcliffe et al., submitted). A cloud-free model was used, along with the priors of Balmer et al. (2025). Given the high spectral resolution and narrow wavelength coverage, the data are expected to probe the upper atmospheric layers above the cloud base, as discussed in Xuan et al. (2022) and Denis et al. (2025). To compare the model to the data, we followed the framework described in Denis et al. (2025).
The results are summarized in Table 1, which reports the planetary RV corrected for barycentric motion, the stellar RV ($12.6\pm 0.3$$\mathrm{km}\,\mathrm{s}^{-1}$; Gontcharov 2006), and the planetary $v\sin i$. The $v\sin i$ measured at the fourth epoch is inconsistent with the other epochs, likely due to instrumental fringing (see Appendix B). We therefore relied on the first three epochs to estimate the rotational broadening and derive $\hbox{$v\sin i$}=9.9\pm 2.3$$\mathrm{km}\,\mathrm{s}^{-1}$, keeping in mind that the fringing effect still affects these data. Assuming $\sin i=1$ and adopting the planetary radius $R_{p}=1.3\pm 0.03,R_{\mathrm{Jup}}$ (Balmer et al. 2025), we can derive an upper limit on the rotation period of the planet ($\sim 16.4\pm 3.8$ h). We thus estimated $P_{\mathrm{rot}}\lesssim 24$ h at 2 $\sigma$.
Table 1: Results on 51 Eri b for the our nights of observations.333
@@ 3.2 Cross-correlation function
To assess the presence of individual molecules in the atmosphere, it is convenient to use a cross-correlation-function (CCF) analysis, in which we correlate molecular templates with the data. In the H band, at the temperature of 51 Eri b, the dominant species are H2O, CH4, and CO. To generate the templates for these molecules, we used the pressure-temperature profiles and the volume-mixing ratio profiles of Exo-REM in tandem with Exo_K. For the derivation of the CCF, we refer the reader to Appendix C.
There appears to be no correlation between the CCF and the derived RV uncertainty. For the first and third epochs, where detection confidence is highest, the derived RV uncertainties are more than twice those of the final epoch. Although it is natural to expect RV uncertainty to correlate with the CCF Signal-to-Noise (S/N), we are in a regime where resolution is also an important parameter. The lower vsini of the last epoch compared to the other epochs keeps the effective resolution of the template higher (see Table 1). This results in sharper variations of the log likelihood near the RV of the planet and therefore less uncertainty in its determination.
Figure 1 shows the combined CCF from the four epochs. Before combining the four CCFs, we shifted each individual CCF to the rest frame of the planet to account for the variations in RV across the different epochs. In addition to the individual molecular templates, we also considered a full template including H2O + CH4 + CO. In general, the molecular detections at individual epochs are marginal, but the combined CCF exhibits a reliable detection of H2O and CH4 and a tentative detection of CO, which confirms the recent findings of Madurowicz et al. (2025).
Figure 1: Combined CCF from the four nights. The combined CCF was computed by shifting the each individual CCF to an RV of 0 $\mathrm{km}\,\mathrm{s}^{-1}$ in order to account for the variations of the RV of the companion across the different epochs.
@@ 4 Updated orbital solution
Table 2: Orbital parameters inferred in this work.444\$a\$\$a\$footnotetext: The periastron passage is computed as $\tau=\frac{t_{\mathrm{p}}-t_{\mathrm{ref}}}{P}$, where tref is the reference time at epoch J2020.0 ($t_{\mathrm{ref}}$ = 58849 MJD) and tp is the time of periastron.
@@
To update the orbit of 51 Eri b, we compiled relative astrometric data from Gemini/GPI and Keck/NIRC2 (Macintosh et al. 2015; Rosa et al. 2020), VLT/SPHERE (Maire et al. 2019), and JWST/NIRCam (Balmer et al. 2025). We used the Markov chain Monte Carlo (MCMC) framework designed by Beust et al. (2016), which makes use of universal Keplerian variables with Stumpff functions (Danby 1987). This formulation achieves more efficient convergence in the case of astrometric orbits sampled over a small orbital arc, as in the present case. The MCMC was run using eight walkers. The burn-in and convergence was assessed using Gelman-Rubin statistics (Gelman et al. 2003; Ford 2006). Before beginning the inference of the chains, we required that $\hat{R}(z)<1.1$ and $\hat{T}(z)>100$ (see Sects. 3.3 and 3.4 of Ford 2006). Then, we let the chains be filled until $\hat{R}(z)<1.01$ and $\hat{T}(z)>1000$. The filled chains constitute the output of the orbital fit.
@@
We did not directly fit for classical orbital elements: the semimajor axis, $a$; the eccentricity, $e$; the inclination, $i$; the argument of periastron, $\omega$; the longitude of the ascending node, $\Omega$; and the time of periastron passage, $t_{p}$. Ford (2006) suggested that working with combinations of these elements allows an optimization of the Markov chains. We specifically fit for $\ln q$, $e\cos(\Omega+\omega)$, $e\sin(\Omega+\omega)$, $sin(i/2)\cos\Omega$, $\sin(i/2)\sin\Omega$, $\ln P$, and $s$. Here, $q$ is the periastron, $P_{q}$ is the orbital period of a circular orbit with radius $q$, and $s$ is the universal variable (a reformulation of the eccentric anomaly; Danby 1987) at a reference epoch. Note that $P$ can also be defined as $P_{q}=2\pi qv_{q}$, where $v_{q}$ is the velocity at periastron. From any solution with those fit parameters, the classical elements can be derived unambiguously.
@@
The updated orbit and associated parameters are presented in Fig. 2 and Table 2. The full posterior distributions of orbital parameters are presented in Fig. 6. The left panel of Fig. 2 shows the predicted RV differences between the planet and the star as a function of time. The middle panel shows a zoomed-in view centered on the HiRISE measurements, and the right panel displays the astrometric orbit of the companion. These plots were generated using 500 orbits randomly drawn from the posterior distributions.
@@
The inclusion of our RV measurements rules out a significant number of orbital solutions, as can be seen in the left panel of Fig. 2. The middle panel shows that a future RV measurement of the planet at the end of 2026 will potentially completely reject the last population of low-eccentricity orbits and place even tighter constraints on the eccentricity. As illustrated in Fig. 6, adding the HiRISE RV measurements significantly improves the orbit. We now estimate the eccentricity to lie within the 0.2–0.7 range at $5\sigma$, with a median at $e=0.55$, completely ruling out very eccentric orbits, above 0.7.
@@
The HiRISE measurements also allowed us to break the typical degeneracy arising with the argument of periastron, $\omega$, and the longitude of ascending node, $\Omega$, when using only astrometric data to fit orbits of exoplanets. This is because the computation of the astrometry of the planet from the orbital parameters is invariant to the transformation ($\omega^{\prime}=\omega+\pi$, $\Omega^{\prime}=\Omega+\pi$), typically producing double-peaked structures in the posterior distributions of these parameters. As a consequence, including our RV measurements enabled the determination of the direction of the motion of the companion along the line of sight. This allowed us to compute a phase curve of the companion (see the right panel of Fig. 7), which is essential for planning follow-up observations using reflected-light imaging or spectroscopy with future instrumentation. This is particularly pertinent in this case, since the reflected-light contrast is predicted to be higher than the thermal-light contrast for this planet, mostly under favorable phase functions (see Appendix. E).